The resting potential of the crab limb nerve is 89 mV. What is the

concentration of potassium ions inside the nerve, if it is 12mmol/1 outside?

Take the temperature equal to 20°C


1
Expert's answer
2022-04-07T13:10:11-0400

The resting potential is given by

"V=60\log\frac{[K^+]_o}{[K^+]_i}\;\rm(mV)"

Hence, the concentration of potassium ions inside the nerve

"[K^+]_i=[K^+]_o\times e^{-\frac{V}{60}}"

"[K^+]_i={12\:\rm mmol/L}\times e^{-\frac{89}{60}}=2.7\:\rm mmol/L"